OpenStudy (anonymous):

How do I Solve by completing the square? m^2-9/2m=3/2

OpenStudy (anonymous):

divide -9/2 by 2 giving you -9/4 square this (-9/4)^2=81/16 add this to both sides giving you m^2-9/2m +81/16 = 3/2 + 81/16 the right hand side is a perfect square (m-9/4)^2=105/16

OpenStudy (anonymous):

sqrt both sides giving you (m-9/4) = +/- sqrt(105)/4

OpenStudy (anonymous):

\[m= \frac{9+/- \sqrt{5}}{4}\]

OpenStudy (anonymous):

remember to add the +/- to any number you square root this is a common mistake people make